# StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models

### v2

@@ 4.4 Models
In line with previous work, we focus our main experiments are based on the API-based LLMs to compare performance fairly. Concretely, ‘gpt-3.5’ level models were used in previous work including current state-of-the-art ActRe (Yang et al., 2024b) and state-of-the-art for in-context learning ReAct (Yao et al., 2023b). Many OpenAI models have become deprecated; notably, all models from ReAct and AdaPlanner Sun et al. (2023) davinci-002, gpt-3.5-turbo-0301, and gpt-3.5-turbo-0613 are now deprecated. Therefore, we re-implemented ReAct and ran the experiments to determine the most suitable model, see Table 1. We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this model for our main experiments.
In our work, we also compare our method using newer state-of-the-art models to show that our method generalises to new models. Specificially, we use the gpt-4o-mini [^4] Again, we chose the smaller model due to cost. and Mixtral (Jiang et al., 2024) models. We note that full GPT-4 level models are prohibitively expensive [^5] A single evaluation run on Alfworld costs approx. \$8 using gpt-3.5 and ReAct, whilst gpt-4 costs 10+ times more.. We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C for the exact settings.
Method Version AW WS
ReAct gpt-3.5-turbo-1106 40.7 18.2
ReAct gpt-3.5-turbo-0125 30.37 14.6
Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS), using available gpt-3.5 models.
4.5 Metrics
In terms of metrics, we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR). Success is a binary metric per each environment in the respective test sets (134 for Alfworld and 500 for Webshop respectively). Success in Alfworld means the agent has successfully completed the whole task. In Webshop, it means the agent has bought an item that has a one hundred percent match with the desired item based on a partially hidden list of shopping item attributes (e.g. the colour, size, price, etc.).
@@ 5 Results
5.1 Alfworld
For Alfworld, we present the results for ReAct, AdaPlanner with and without code execution, and StateAct (ours), which consists of goal + state + thought + action. We also show StateAct without each of the components (i.e. without goal, state and thought). Interestingly, we find that, contrary to previous findings, ‘thought’ or ‘reasoning’ actually sometimes hinders the performance.
In Table 2, we can see that StateAct with all goal + state + thought and the correction performs the best. It outperforms ReAct with correction by around 13 points (when using the same GPT model for ReAct) and by around 9 points (when using the better model for ReAct). StateAct also outperforms ReAct by 22 points when corrections are not used. Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command-mapping (similar to our correction) along with using code-execution.
Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16- and 23-point jump for ReAct and a 27-point jump for StateAct. This indicates that the model generally performs very well; however, it struggles with minute differences in domain-specific syntax.
Method GPT-3.5 SR %
Baselines w/o corrections
ReAct 0301* 51.9
ReAct 1106 40.7
ReAct 0125 30.37
ReAct (joined***) 1106 36.30
Results w/o corrections
StateAct (ours) 1106 50.37
- w/o thought 1106 62.96
- w/o goal 1106 42.96
- w/o state 1106 44.44
Baselines with corrections
ReAct + corr. 0301* 68.15
ReAct + corr. 1106 63.70
AdaPlanner No-Exec** 0301* 46.66
AdaPlanner 0301* 75.56
Results with corrections
StateAct + corr. (ours) 1106 77.04
- w/o thought 1106 64.44
- w/o goal 1106 74.04
- w/o state 1106 64.44
Table 2: Success Rate (SR) on the 134 test-set examples from Alfworld. *gpt-3.5-0301 was deprecated in June 2024. **No-Exec means AdaPlanner without code execution. ***joined means that thought + action are produced at every turn. AdaPlanner results are from Sun et al. (2023). All other results are single run.
5.2 Webshop
For Webshop, we present the results for ReAct and StateAct. Similarly to Alfworld, we also present the results of StateAct without each of goal, state and thought, see Table 3. Interestingly, we see that removing thought produces the highest results and outperforms ReAct by 10 points. Our hypothesis is that domain-specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
Method GPT-3.5 Success Rate %
@@ 5.2 Webshop
Rule-based* N/A 9.60
RL* N/A 17.60
ReAct 1106 17.80
Results
StateAct (ours) 1106 17.00
- w/o thought 1106 27.80
- w/o goal 1106 20.40
- w/o state 1106 21.00
Table 3: Success Rate (SR) on the 500 test-set examples from Webshop. *results taken from Yao et al. (2023a). For StateAct without thought and ReAct Results are the average of two runs. Other Results are single run.
5.3 Results using additional models
We validate our results further with additional LLMs. Again we see a significant boost in performance, see Table 4. We experimented with gpt-4o-mini and Mixtral [^6] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’. (Jiang et al., 2024) on Alfworld. ReAct achieves 68.15 and 72.59 respectively, while StateAct achieves 71.85(+3.7) and 83.70(+11.2).
Method Model SR %
Baselines
ReAct Gpt-4o-mini 68.15
ReAct Mixtral-8x22B* 72.59
Results
StateAct (ours) Gpt-4o-mini 71.85
StateAct (ours) Mixtral-8x22B* 83.70
Table 4: Success Rate (SR) on the 134 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtral is ‘mixtral-8x22b-instruct-v0.1’.
5.4 Summary of results
In conclusion, we found that our simple goal-reminding and state-tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more than 10 points and even outperforms leading approaches that rely on code-execution. Interestingly, we found that the approaches are quite sensitive to domain-specific syntax and, when this is the case, ‘thoughts’ that are verbose can hinder performance.
6 Analysis and Ablations
In the results section, we discovered that our methods perform better than previous state-of-the-art. Our method shows strong performance with in-context learning without resorting to additional tools, data or bigger models. In this section, we analyse our results further and also answer our second hypothesis that goal ‘reminding’ and state-tracking help with long-range reasoning. For all ablation studies, we focus on Alfworld as it has two favourable properties over Webshop. Firstly, Alfworld has a longer time horizon (50 steps vs. 15 in Webshop), with tasks taking, on average, less than 10 steps in Webshop and around 20 to 30 steps in Alfworld. Secondly, Alfworld has much less domain-specific syntax and is purely text-based, while Webshop has a more specific syntax to follow.
6.1 Do goal reminders help with long-range tasks?
For this purpose, we compare the original ReAct (thought + action) with the goal included, i.e. StateAct (goal + thought + action). In Figure 5 we can see that, while the performance of both ReAct and StateAct goes down as there are more steps, the goal tracking has a substantially better relative performance as the number of steps increases. This finding is in-line with our original motivation that LLM agents deteriorate in performance as the prompts and interactions get longer.
To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct [^7] We ignore ‘thought’ turns for ReAct as otherwise ReAct would have even more steps. and StateAct. Table 5 clearly show that ReAct, with an average of 38.84 steps to solve an environment, is the least efficient whilst StateAct, with an average of 28.96 steps to solve an environment, is the most efficient. This shows that not only does goal reminders help with longer range tasks, it also helps with efficiency, by shortening the tasks. See Appendix E for more discussion.
Figure 5: Goal vs. No Goal, on the test set of Alfworld, using gpt-3.5-turbo-1106 without correction.
6.2 What effects does state-tracking have?
We also analyse whether state-tracking helps with long-range reasoning and efficiency. We compare the full StateAct against StateAct without state-tracking, as well as comparing ReAct (thought + action) against StateAct with state-tracking (state + thought + action). In Figure 6 we see that state-tracking also helps with long-range reasoning. In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with both state tracking and goal-tracking longer-range tasks can be solved. Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^8] In terms of cost, we found that, despite our method using a twice-longer prompt, our cost remains similar, at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer steps.. Concretely, ReAct has 24/134 examples that are in the bucket ‘40-50’ and solves 0, while our method has 6/134 examples and solves 4.
Therefore, we find that explicit state-tracking helps with long-range tasks and enables the agent to solve the tasks more efficiently.
Figure 6: State vs. No State, on the 134 test examples from Alfworld. gpt-3.5-turbo-1106 with correction.
Method Avg. Steps $\downarrow$
ReAct 38.84
StateAct (goal+thought+action) 31.19
StateAct (full) 28.96
ReAct with corr. 31.49
StateAct with corr. 19.11
- w/o thought 23.76
- w/o goal 20.09
Table 5: Average number of Steps (Avg. Steps) [lower is better] on test-set of Alfworld, using gpt-3.5-1106.
6.3 Does the model perform actual state-tracking?
We ask ourselves whether the model is actually performing state-tracking. For that purpose, we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^9] See Appendix D for details. based on the actions the agent takes. For example, if the agent produces the action go to fridge 1 and the environment accepts this action, we update the state with current location: fridge 1. We compare the ‘gold’ state against the predicted state. Figure 7 shows that StateAct does in fact correct state-tracking 88% of the time. We also observe that thoughts and goals help the state-tracking.
6.4 Does json structure help with performance?
Since we found that domain-specific syntax hinders performance, we investigated whether adding a structured format like json would help. For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix G for more details. Surprisingly, we found that the json format hinders performance significantly, see Table 6. However, we also see that corrections help the json format less, indicating that json helps with syntax but hinders performance.
Method SRSR (+json)%
StateAct 50.37 45.19 (-5.2)
StateAct w/ corr. 77.04 58.52(-18.5)
7 Conclusion and future work
Inspired by the challenge that LLMs and LLM agents struggle with long context, we propose a novel in context learning method for LLM agents - StateAct.
We introduce ‘chain-of-states’, which relies on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution. StateAct outperforms the previous state-of-the-art that uses in-context learning alone, by between 9% and 20% given different models and tasks, and outperforms in-context learning with tools (i.e. code-execution) by 3%, while performing on par with methods that require additional training data.
In conclusion, we show that explicit state-tracking and goal reminders help the model solve long range reasoning tasks and interestingly also make the model more efficient.
@@ 8.1 Computational footprint
Running many of the experiments presented in this paper can have a significant computational footprint. We should consider the environment and financial resources for reproduciblity of our work. We aimed to address this concern by using gpt-3.5-turbo level models, reporting costs and minimising the cost of our method.
@@ 9.1 Languages and evaluation benchmarks
We evaluated our method only in the English language and on two evaluation benchmarks. While we do not expect major changes in other languages, this is something that should be investigated. Furthermore, performance on other benchmarks should be evaluated as well.
@@ B.3 License
Appendix C Code snippet to call OpenAI / GPT-3.5

### v3

@@ 4.5 Models
In our work, we compare our method using newer state-of-the-art models, architectures and sizes to show that our method generalises. Specifically, we use Mistral-Small-24B-Instruct (Jiang et al., 2023), Qwen-2.5-7B,14B,32B-Instruct (Qwen et al., 2025) and Gemma2-27B-Instruct (Team et al., 2024). We note that api-based models such as OpenAI’s models are generally very expensive [^2] A single evaluation run on Alfworld costs approx. \$8 using gpt-3.5 and ReAct, whilst gpt-4 costs 10+ times more. and lack rigorous reproducibility standards [^3] Since we do not have access to weights and inference settings and models become regularly deprecated.. Nonetheless, we include experiments using gpt-3.5 and gpt-4o-mini. We use temperature 0 for all experiments and sample only the top 1 response; we use vllm for inference (Kwon et al., 2023), see Appendix D & E for details.
@@ 5 Results
Overall we present results of state-of-the-art methods such as ActRe and ADaPT, as well as the underlying base agent ReAct. In Table 1 we see that while methods that rely on ReAct + extensions, such as ActRe and AutoGuide outperform StateAct overall, they also rely on additional training data and computation. Furthermore, StateAct achieves the best result among base agents, outperforming ReAct between 7% and 30% and achieves comparable result with the state-of-the-art methods, while not using any additional tools or data. For example, AutoGuide (ReAct + RAG) gets 0.79, while StateAct gets 0.77.
Method Score
State-of-the-art methods
AdaPlanner (Code-prompt + exec.)1 0.75
AutoGuide (ReAct + RAG)2 0.79
ActRe (ReAct + fine-tuning)3 0.83
ADaPT (ReAct + test time scaling)4 0.72
Base agents
Act (few-shot only) 0.41
ReAct (few-shot only) 0.64
AdaPlanner (Code-prompt only)1 0.45
StateAct (ours, few-shot only) 0.77
Table 1: Results on the 134 test samples of Alfworld using gpt-3.5. ReAct and StateAct scores are single run with greedy decoding and gpt-3.5-1106. 1=code-execution (Sun et al., 2023), 2=(Fu et al., 2024), 3=(Yang et al., 2024), 4=(Prasad et al., 2024).
5.1 Base agent comparison
Agent Name Mistral-24B Qwen-7B Qwen-14B Qwen-32B Gemma-27B Average $\uparrow$
Alfworld*
ReAct 0.44 0.10 0.75 0.89 0.71 0.58
StateAct 0.49 0.46 0.78 0.90 0.68 0.66
Webshop**
ReAct 0.34 0.19 0.22 0.27 0.26 0.26
StateAct 0.35 0.12 0.33 0.32 0.29 0.28
Textcraft***
ReAct 0.33 0.02 0.31 0.31 0.18 0.23
StateAct 0.40 0.04 0.37 0.40 0.34 0.31
Table 2: Base agent performance across different models and environments. *=134 Test Environments from Alfworld. **=100 Test Environments from Webshop. ***=100 Test Environments from Textcraft. M=Mistral-Instruct-2501, Q=Qwen2.5-Instruct, G=Gemma 2-Instruct. We use greedy decoding (temperature=0).
Since ReAct is the previous best base agent and forms the basis of state-of-the-art approaches, we compare against it in detail. The results in Table 2 show that StateAct outperforms ReAct across three different benchmarks and five different models. Sometimes the difference is quite substantial, with StateAct outperforming ReAct by more than 10 points. Across all 15 experiments only in two ReAct performs better than StateAct: in Webshop, where the 7B model is likely overwhelmed by the amount of textual input it receives, as Webshop has a lot of verbal input; and in the case of Gemma2-27B ReAct slightly outperforms StateAct on Alfworld; a hypothesis is that Gemma has a limited context length of 8192, while Alfworld requires long traces due to the longer step length.
We further validate our results with additional LLMs that are too large to fit on a single GPU or are closed source. In Table 3 we see a significant performance increase. Using gpt-3.5, gpt-4o-mini and Mixtral-8x22B [^4] Mixtral-8x22b-instruct-v0.1 was queried using Nvidia’s NIM API https://developer.nvidia.com/nim [Last Accessed March 2025]. (Jiang et al., 2024) on Alfworld, ReAct achieves 63.7, 68.15 and 72.59, while StateAct achieves 77.04 (+13.3), 71.85(+3.7) and 83.70(+11.2) respectively.
Method Model Success Rate %
ReAct Gpt-3.5 0.64
StateAct Gpt-3.5 0.77
ReAct Gpt-4o-mini 0.68
StateAct Gpt-4o-mini 0.72
ReAct Mixtral-8x22B 0.73
StateAct Mixtral-8x22B 0.84
Table 3: Results on the 134 test examples from Alfworld. Results are single run and greedy. Models used: gpt-3.5-1106, gpt-4o-2024-07-18, mixtral-8x22b-instruct-v0.1.
5.2 Base agents + test time scaling
An important contribution is to validate that StateAct can be used as a drop-in replacement with advanced methods. To this end we validate StateAct using test-time scaling using the ADaPT method. Starting with ADaPT, which is based on ReAct, as the starting point(Prasad et al., 2024), we enhance their method using StateAct. In Table 4 we can clearly see that StateAct scales well with test time scaling jumping in performance by 39%, and surpassing ADaPT+ReAct by 12%.
Model/Agent Name Normal +ADaPT
Mistral-24B
ReAct 0.33 0.53
StateAct 0.40 0.64
Qwen2.5-7B
ReAct 0.02 0.09
StateAct 0.04 0.11
Qwen2.5-14B
ReAct 0.31 0.55
StateAct 0.37 0.53
Qwen2.5-32B
ReAct 0.31 0.64
StateAct 0.40 0.62
Gemma-2-27B
ReAct 0.18 0.19
StateAct 0.34 0.35
Average
ReAct 0.23 0.4
StateAct 0.31 0.45
Table 4: Comparison of test-time scaling performance on 100 test samples from Textcraft. Normal refers to using just the base agent. +ADaPT means running the respective base agent with test time scaling using the ADaPT method. ADaPT code is adapted to run ReAct and StateAct. We use greedy decoding and $d_{max}$ = 2.
5.3 Summary of results
StateAct establishes itself as the best-performing base agent, surpassing ReAct by 7–30% across multiple benchmarks while requiring no additional training data or external tools. Although advanced methods like ActRe and AutoGuide achieve higher scores, they rely on costly training and retrieval. We also validate StateAct as a drop-in replacement for ReAct in test-time scaling. These results highlight the performance and usability of StateAct as a robust foundation for LLM-based agents.
6 Analysis and Ablations
In the results section, we discovered that our methods outperform the previous state-of-the-art base agent, ReAct. Our method shows strong performance with in-context learning without resorting to additional tools or data. In this section, we analyse our results further and also show that self prompting and state-tracking help with long-range reasoning. For most ablation studies, we focus on Alfworld as it has two favourable properties over Webshop and Textcraft. Firstly, Alfworld has a longer time horizon (50 steps vs. 15 in Webshop, 40 in Textcraft). Secondly, Alfworld is more realistic than Textcraft, as Alfworld is a robotics environment, while Textcraft is based on a game. For completeness, we include ablation results in Table 6 and full results in Appendix L,M,N.
6.1 Does self-prompting help with long-range tasks?
For this purpose, we compare the original ReAct (thought + action) with the self-prompting included, i.e. StateAct (goal + thought + action). In Figure 5 we can see that, while the performance of both ReAct and StateAct goes down as there are more steps, the goal tracking has better relative performance as the number of steps increases and is able to solve longer tasks of 40 to 50 steps. This finding is in line with our original motivation, that LLM agents deteriorate in performance as the prompts and interactions get longer.
To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct [^5] We ignore ‘thought’ turns for ReAct as otherwise ReAct would have even more steps. and StateAct. Table 5 clearly shows that ReAct, with an average of 31.49 steps to solve an environment, is the least efficient whilst StateAct, with an average of 19.11 steps to solve an environment, is the most efficient. This shows that not only does self-prompting help with longer range tasks, it also help with efficiency, by shortening the tasks. See Appendix G for more discussion.
6.2 What effects does state-tracking have?
We also analyse whether state-tracking helps with long-range reasoning and efficiency. We compare the full StateAct against StateAct without state-tracking, as well as comparing ReAct (thought + action) against StateAct with state-tracking (state + thought + action). In Figure 5 we see that state-tracking also helps with long-range reasoning. In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with both state tracking and goal-tracking longer-range tasks can be solved. Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^6] Despite StateAct using a twice-longer prompt, our cost remains similar to ReAct, at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer steps.. Concretely, ReAct has 24/134 examples that are in the bucket ‘40-50’ and solves 0, while StateAct has 6/134 examples and solves 4. Therefore, we find that explicit state-tracking helps with long-range tasks and to solve the tasks more efficiently.
Figure 5: State vs. No State, on the 134 test examples from Alfworld, using gpt-3.5-turbo-1106
Method Avg. Steps $\downarrow$
ReAct 31.49
StateAct (goal, state, thought) 19.11
- w/o goal (self-prompt) 20.09
- w/o state 22.50
- w/o thought 23.76
Table 5: Average number of steps (Avg. Steps) [lower is better] on the test set of Alfworld, using gpt-3.5-1106.
6.3 Does the model perform actual state-tracking?
We investigate whether the model is actually performing state-tracking. For that purpose, we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^7] See Appendix F for details. based on the actions the agent takes. For example, if the agent produces the action go to fridge 1 and the environment accepts this action, we update the state with current location: fridge 1. We compare the ‘gold’ state against the predicted state. Figure 6 shows that StateAct achieves a state-tracking accuracy of 88%. We also observe that thoughts and goals help the state-tracking.
6.4 Do ‘thoughts’ help?
While chain-of-thought helps in most cases, surprisingly, we see that in some cases chain-of-thought can be harmful. Specifically, in Webshop ‘thoughts’ actually harm overall performance across different agent and model types, see Table 6. A hypothesis is the verbosity of the Webshop environment, thus confusing the model with ‘verbal’ thoughts.
Agent Name AW WS TC
@@ 6.4 Do ‘thoughts’ help?
Act 0.51 0.19 0.27
Thought+Act (ReAct) 0.58 0.26 0.23
Our Methods
State+Act 0.51 0.31 0.22
State+Thought + Act 0.58 0.18 0.34
Goal+Act 0.63 0.24 0.26
Goal+Thought + Act 0.65 0.18 0.30
Goal+State+Act 0.56 0.29 0.21
Goal+State+Thought+Act 0.66 0.28 0.31
Table 6: Ablation table. Averaged results across Mistral, Qwen, Gemma on AW=Alfworld, WS=Webshop and TC=Textcraft. Goal=Self-prompt.
7 Conclusion
Our work is driven by the fundamental challenge that LLM agents struggle with long context and keeping on track with instructions. The current state-of-the-art to overcome such challenges propose extensions on top of the base agent, ReAct. In contrast, we introduced a novel base agent, StateAct, an in-context learning method that leverages chain-of-states and self-prompting to significantly enhance the capabilities of LLM agents. By rethinking how agents track and utilize state information, StateAct establishes a new state-of-the-art for base agents, surpassing ReAct by 9% to 20% across different models and tasks. Furthermore, we demonstrate that StateAct scales effectively when combined with advanced techniques such as ADaPT, reinforcing its usability.
Beyond raw performance, our analysis uncovers a crucial insight: StateAct not only improves reasoning but also enhances efficiency, allowing agents to achieve better results with fewer steps. This suggests that integrating structured state-tracking and self-prompting cues helps mitigate the well-documented long-context issue in LLMs.
Our findings suggest that methods like StateAct can serve as a practical and efficient way to improve LLM-based agents without requiring costly external augmentations or data such as retrieval or fine-tuning. By enabling agents to manage their own state explicitly, StateAct provides a scalable approach to improving reasoning and decision-making across a range of tasks. This makes StateAct a viable drop-in solution for current systems, offering both better performance and greater efficiency in LLM agent tasks.
@@ 8.1 Computational footprint
Running many of the experiments presented in this paper can have a significant computational footprint. We should consider the environment and financial resources for reproduciblity of our work. We aimed to address this concern by models that are less computationally demanding such as gpt-3.5-turbo level models or open source models that fit on a single GPU (A100, 80GB), reporting costs and minimising the cost of our method.
@@ 9.1 Languages and evaluation benchmarks
We evaluated our method only in the English language and on three evaluation benchmarks. While we do not expect major changes in other languages, this is something that should be investigated.
@@ A.2 Action Types
A.2.1 Alfworld correction
Note: In the latest release of Alfworld (December 2024) this was fixed by replaceing the put <obj> in/on <place> command with move <obj> to <place> command. In our work we report results on the latest version of alfworld.
@@ B.3 License
Appendix C Textcraft
C.1 Commands and environment
Textcraft has one environment type: ‘craft’, as well as three commands: 1. inventory, 2. craft, 3. get
C.2 Crafting Recipes
Textcraft has crafting recipes that range from easy to hard. Where hardness is measured by the ‘depth’ of the crating recipe. Specifically, depths of 2, 3 and 4 are present in the dataset.
C.3 License
Textcraft is published under the permissible MIT license.
Appendix D Compute Requirements for local LLMs
The exact code will be released upon publication. However, to help reproducibility we ran all experiments on single A100 80GB GPUs. In terms of software we used: vLLM for inference. The hyper-parameters were set to: max model length 16000 (except for Gemini, where we used 8192), temperature = 0, datatype="auto" (which results in bfloat16).
Appendix E Code
E.1 Code Snippet to call Local LLMs
from vllm import LLM, SamplingParams
self.llm = LLM(
model=model,
tensor_parallel_size=tensor_parallel_size,
gpu_memory_utilization=0.95,
max_model_len=max_model_len,
dtype="auto"
messages = [
# {"role": "system", "content": "You are a helpful assistant."},
{"role": "user", "content": prompt[-self.max_model_len:]}
text = self.tokenizer.apply_chat_template(
messages,
tokenize=False,
add_generation_prompt=True
sampling_params = SamplingParams(
temperature=self.temperature,
top_p=1.0,
repetition_penalty=1.00,
max_tokens=min(2000,self.max_model_len),
stop = self.stop_sequences,
seed = self.seed
outputs = self.llm.generate([text], sampling_params)
return outputs[0].outputs[0].text
E.2 Code Snippet to call OpenAI / GPT-3.5